# Rectified Flows for Fast Multiscale Fluid Flow Modeling

### v1

@@ 4.1 Mathematical Formulation
When $S(v)$ is small, the velocity field $v(\cdot,\tau)$ remains nearly constant in $\tau$, ensuring trajectories are close to straight lines in state space.
ODE formulation. Sampling from the rectified flow amounts to solving the ODE,
$\frac{du_{\tau}}{d\tau}\;=\;v\bigl{(}u_{\tau},\,\tau\bigr{)},\quad u_{0}\sim\mu_{0}.$ (4.2)
As $\tau$ goes from 00 to $1$, $u_{\tau}$ travels along straighter interpolants between $\mu_{0}$ and $\mu_{1}$. In practice, we learn a $v$ that approximates the barycentric velocity field while promoting small $S(v)$.
Practical learning objective. Given samples $\{(u_{0},u_{1})\}$ from $\gamma$, one can minimize
@@
which precisely encourages small $S(v)$. This ensures $v$ approximates the conditional expectation $u_{1}-u_{0}$ at each partial interpolation $U_{\tau}$, but remains as “straight” as possible over $\tau$. Furthermore, by integrating over these noise scales, the model sees high frequencies more consistently than an MSE-based approach.
Conditional expectation viewpoint. Equivalently, if $(U_{0},U_{1})\sim\gamma$ and $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$, then the true minimizer takes the form
$v(u,\tau)\;=\;\mathbb{E}[\,U_{1}-U_{0}\,\mid\,U_{\tau}=u\,].$
by the $L^{2}$ optimality property of conditional expectations. Thus $v(u,\tau)$ is the average instantaneous displacement from $u_{0}$ to $u_{1}$, conditioned on arriving at $u$ at time $\tau$.
@@ 4.2 Implementation Details for PDEs
1:Dataset $\{(u_{i},u_{f})\}$ with $u_{f}=\mathcal{S}(u_{0})$; velocity model $v_{\theta}$; noise schedule $\sigma(\tau)$
@@
where $u_{\tau}$ is a low-frequency representation of $u_{f}$, $u_{i}$ is the initial datum, and $\Gamma(\cdot)$ is an MLP-based time embedding:
@@
Within the U-Net (see SM A.1), we alternate residual blocks with multi-head attention layers, downsampling to a bottleneck and then upsampling via skip connections. A final $1\times 1$ projection recovers either the field increment $\hat{u}\approx u_{1}-u_{0}$ or a noise residual. By injecting the rectification time $\Gamma(t)$ into each block, the model performs a fully deterministic trajectory integration instead of stochastic diffusion.
@@
$\frac{du_{\tau}}{d\tau}=v_{\theta}\bigl{(}u_{\tau},\,u_{i},\,\tau\bigr{)},\quad u_{\tau}|_{\tau=0}=\xi,\quad\xi\sim\mathcal{N}(0,I).$
@@
Efficiency and accuracy. Because rectified flows explicitly encourage a “straighter” velocity field, accurate solutions often require fewer integration steps than their diffusion-based counterparts. This advantage can be crucial in computational fluid dynamics and other PDE applications where each solver step is expensive. See SM D.3 and F for a detailed comparison of step-count vs. accuracy trade-offs.
@@ 4.3 Computational Efficiency Comparison
Figure 4.1: Inference cost versus model size. Rectified flows achieve the desired accuracy with only 44448888 ODE steps, while GenCFD variants need 64646464128128128128 reverse-time SDE steps. We vary U-Net depth from $4$ to $8$ residual blocks, growing the parameter count from $5.5$ M to $10.2$ M.
Diffusion surrogates (e.g. GenCFD) generate samples by integrating a reverse-time SDE whose trajectory bends through latent noise; even with accelerated samplers such as DDPM or DPM-solver the procedure still requires tens of score-network evaluations. Rectified flows, in contrast, follow a single deterministic ODE whose nearly straight path tolerates large or adaptive steps, so inference needs only a handful of evaluations. Figure 4.1 quantifies this gap, showing how the runtime of diffusion models scales almost linearly with the number of steps, whereas rectified flows remain fast even as network size grows.
@@ 5.1 Experimental Setup
We benchmark our method, termed henceforth as ReFlow, against four neural surrogate models. GenCFD and GenCFD $\circ$ FNO are conditional diffusion surrogates following the framework introduced earlier in this section, whereas FNO and UViT serve as one-shot deterministic baselines. Detailed descriptions of all baselines are provided in SM C.
@@
    Shear Layer in 2D. This setup models the evolution of an incompressible shear layer, initially concentrated near a horizontal interface, which rolls up into intricate vortical structures over time. We generate approximate solutions by numerically solving the 2D Navier–Stokes system (with periodic boundaries) at a resolution of $512^{2}$, then downsample to $128^{2}$ for training and testing. The data has been generated with the spectral viscosity code [30, 27].
@@
    Cloud-Shock Interaction in 2D. In this popular compressible-flow test, a shock wave moves through a higher-density “cloud” region, triggering complex wave interactions and small-scale turbulent mixing in the wake. We solve the compressible Euler system at an effective resolution of $256^{2}$ using a GPU-optimized finite-volume ALSVINN code of [19, 6], then also downsample to $128^{2}$.
@@
    Richtmeyer–Meshkov in 2D. In this case, a shock wave interacts with a density interface separating two fluids, depositing vorticity and triggering vigorous mixing. We numerically solve the compressible Euler system on a $256^{2}$ grid, then downsample to $128^{2}$ to create the training and test data. The data has been generated with a high-resolution finite-volume ALSVINN code of [6]
@@
being the probability distribution specified in the Supplementary Material (SMB.2). For each $u_{0}$, the PDE solver (e.g. a spectral or finite-volume code) is run forward in time to produce high-fidelity snapshots, which serve as training data.
@@
To create a “ground truth” ensemble for comparison under a fixed $\widehat{u}_{0}$, we employ an ensemble perturbation procedure (detailed in SM  B.2), where $\widehat{u}_{0}$ is slightly perturbed to generate multiple nearby initial states. Each perturbed state is then advanced using a high-resolution PDE solver (e.g. the Azeban code for incompressible flows, or a finite-volume code for compressible problems), yielding a set of flow realizations. This set approximates the conditional measure in question, i.e. $p\!\bigl{(}u(t)\mid\widehat{u}_{0}\bigr{)}$.
@@
    Distribution shift. We consider Dirac initial data $\delta_{\widehat{u}_{0}}$ where $\widehat{u}_{0}\sim\nu$, for a different distribution $\nu\neq\mu$. This scenario probes the ability of our model and baselines to generalize under shifts in the initial condition distribution. Specific details on $\nu$ are provided in SM B.2.
We will show results in both regimes to assess how well rectified flows can capture the spread of chaotic solutions starting from fixed initial states—and how it compares to baselines.
@@ 5.2 Evaluation Metrics
We evaluate the generative performance of each method using three main metrics, with details provided in SM D:
@@
Table 5.1: Comparison of models across datasets CS2D, SL2D, and RM2D, in terms of mean relative error ($e_{\mu}$), standard deviation relative error ($e_{\sigma}$), and Wasserstein-1 distance ($W_{1}$). Best-performing model is colored in Blue and second-best performing model in Orange.
@@
Baselines and Evaluation Settings. To stay within reasonable training times for our infrastructure, we restrict the total number of training iterations up to 120,000 for each model and recorded the best achievable results. Under this budget, only a reduced-capacity GenCFD (with two downsampling stages in its U-Net backbone, versus three in Rectified Flow) can match the accuracy reported in Table 5.1. We also evaluated larger GenCFD variants trained for 500,000 iterations, but observed no meaningful improvement over the results shown here. Meanwhile, at inference time Rectified Flow requires only 8 denoising steps to achieve comparable or superior accuracy, whereas diffusion baselines (including GenCFD) need 128 steps, underscoring RF’s efficiency in both training and sampling. FNO and UViT also require smaller models than ReFlow for optimal performance.
Since FNO, and UViT produce deterministic predictions, we approximate their spread by an ensemble perturbation strategy (SM  B.2), where each baseline is fed slightly perturbed inputs and outputs are aggregated to form a sample set. In contrast, both the original GenCFD (baseline) and our ReFlow are inherently probabilistic, requiring no auxiliary procedure to generate diverse samples.
Figure 5.1: Energy-spectrum evolution for a random cloud–shock initial datum. ReFlow (blue) tracks high-wavenumber energy faster and more accurately than GenCFD (red).
@@
Moreover as visualized in the SM, ReFlow generates samples of very high quality, matching the qualitative features of the ground truth and capturing features such as (interacting) vortices as well as propagating shock waves very accurately. Another aspect where ReFlow really shines in being able to generate the correct point pdfs. Although already indicated in the very low 1-point Wasserstein distance errors from Table 5.1, we present 1-point pdfs in the SM to illustrate this observation.

### v2

@@ 4.3.3 Capacity–coverage–fit decomposition for one-step RF
Proposition 4.12 (One-step law-level error: capacity, coverage, and fit).
Under the structure-function modulus assumption (17), the one-step $W_{2}$ error between the PDE law and the RF law satisfies
$W_{2}\!\bigl(\mu_{t+\Delta t},\,\widehat{\mu}_{t+\Delta t}\bigr)\;\leq\;\Bigl(C_{d}\,\omega\!\bigl(c_{d}/K_{c}\bigr)\;+\;\varepsilon_{\mathrm{train}}^{2}(t)\Bigr)^{1/2},$ (22)
where the first term is a coverage error determined by the structure-function modulus and the cutoff $K_{c}$, and the second is a fit error determined by training.
Proof sketch.
Couple $u\sim\mu_{t}$ and $\widehat{u}$ using the same input; then $W_{2}^{2}(\mu_{t+\Delta t},\widehat{\mu}_{t+\Delta t})\leq\mathbb{E}\|S_{\Delta t}u-\widehat{u}\|_{L^{2}}^{2}$. Insert $P_{\leq K_{c}}S_{\Delta t}u$, apply the triangle inequality, and bound the tail term by Lemma 4.9 (applied at time $t+\Delta t$). ∎
Interpretation and experimental validation.
At a fixed resolution (fixed $K_{c}$), Proposition 4.12 says that even a perfectly trained rectified flow cannot beat the coverage term $\omega(c_{d}/K_{c})^{1/2}$ dictated by small-scale physics. This is precisely why spectrum matching (Figure 13) is a non-negotiable diagnostic on multiscale flows: if the surrogate loses the high-frequency tail, it is necessarily paying a coverage penalty in law. The remaining error is controlled by the training fit and the straightness properties of the learned velocity (via the continuous-time bounds in Section 4.1). Rectified flows improve efficiency precisely by reducing the curvature term in these bounds, so that a given coverage+fit accuracy can be reached with many fewer ODE steps than a diffusion-based surrogate such as GenCFD (Section 5.6).
5 Correctness and Curvature-Controlled Discretization
Organization.
This section states the main correctness and discretization results that underpin the narrative above. Full proofs are deferred to the theory appendix (to be written separately); we keep only the steps needed to make the dependence on fit, straightness, and discretization explicit in the main text.
5.1 Exact transport under the barycentric rectified velocity
Proposition 5.1 (Exact transport under the barycentric velocity).
Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathcal{U})$ and let $\gamma\in\mathcal{P}(\mathcal{U}\times\mathcal{U})$ be a coupling of $(\mu_{0},\mu_{1})$. Define $T_{\tau}(u_{0},u_{1})=(1-\tau)u_{0}+\tau u_{1}$ and $\rho_{\tau}:=(T_{\tau})_{\#}\gamma$. Let $v_{\star}(\cdot,\tau)$ be the barycentric velocity field of defined in 2. Assume (e.g. on a finite-dimensional Galerkin discretization of $\mathcal{U}$) that the continuity equation
$\partial_{\tau}\rho_{\tau}+\nabla\!\cdot(\rho_{\tau}\,v_{\star}(\cdot,\tau))=0$ (23)
admits a unique regular Lagrangian flow $X_{\tau}$ associated to $\dot{u}_{\tau}=v_{\star}(u_{\tau},\tau)$. Then $(X_{1})_{\#}\mu_{0}=\mu_{1}$.
Proof sketch.
By construction, $\rho_{\tau}$ is the law of $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$ for $(U_{0},U_{1})\sim\gamma$. Differentiate $\mathbb{E}[\varphi(U_{\tau})]$ and condition on $U_{\tau}$ to identify the weak form of (23). Uniqueness of the continuity equation solution (in the chosen Galerkin setting) implies $\rho_{\tau}=(X_{\tau})_{\#}\rho_{0}$, hence $\rho_{1}=(X_{1})_{\#}\rho_{0}$, i.e. $\mu_{1}=(X_{1})_{\#}\mu_{0}$. ∎
Why this matters. Proposition 5.1 isolates a clean correctness statement: if $v_{\theta}$ matches the barycentric rectified velocity $v_{\star}$ sufficiently well (in a sense quantified by Theorem 4.2), the induced flow pushes forward $\mu_{0}$ to $\mu_{1}$ at the level of laws.
5.2 Curvature controls the required number of ODE steps
Corollary 5.2 (Global Euler error controlled by time-curvature).
Assume the hypotheses of Lemma 4.4 for $\dot{u}=v(u,\tau)$ on $\tau\in[0,1]$ and assume $v(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$. Let $u^{\Delta}$ be the explicit-Euler approximation with uniform step $\Delta\tau=1/N$. Then there exists a constant $C=C(L)$ such that
$\sup_{\tau\in[0,1]}\|u^{\Delta}_{\tau}-u_{\tau}\|\;\leq\;C\,\Delta\tau\,\int_{0}^{1}\Big(\|\partial_{\tau}v(u_{\tau},\tau)\|+\|J_{u}v(u_{\tau},\tau)\,v(u_{\tau},\tau)\|\Big)\,d\tau.$ (24)
In particular, when $\int_{0}^{1}\|\partial_{\tau}v(u_{\tau},\tau)\|\,d\tau$ is small (“straight” flow), the $N$ required to reach a fixed tolerance is proportionally smaller.
Proof sketch.
Sum the local truncation errors from Lemma 4.4 over $N$ steps and propagate them using stability of the flow under the $L$-Lipschitz condition via Grönwall. ∎
Connection to our curvature proxy. The EMA deviation $s_{t}=\|v_{t}-v_{t}^{\mathrm{ema}}\|$ used in (6)–(7) is a computable proxy for the same time-variation that drives (24). Blending reduces effective curvature; step control reduces $\Delta\tau$ when curvature spikes. This directly rationalizes why ReFlow can reach the “spectrum-correct” regime in very few steps (Figure 13), and why the same accuracy level is reached at $8$ steps (Table 1) whereas diffusion baselines typically need $\geq 128$.
5.3 A compact regularity interpretation of structure functions
Lemma 5.3 (Structure functions imply Besov regularity (heuristic equivalence)).
Let $u\in L^{2}(\mathbb{T}^{d})$ and define the increment seminorm
$u]_{B^{s}_{2,\infty}}:=\sup_{0<|h|\leq 1}\frac{\|u(\cdot+h)-u(\cdot)\|_{L^{2}}}{|h|^{s}}.$
If the second-order structure function satisfies $S_{r}^{2}(u)\lesssim r^{2s}$ for all $r\in(0,1]$, then $u]_{B^{s}_{2,\infty}}\lesssim 1$. Consequently, a law-level modulus bound $S_{r}^{2}(\mu)\lesssim r^{2s}$ implies that typical samples from $\mu$ have Besov smoothness $s$, consistent with the Fourier-tail decay in Lemma 4.9.
Remark 5.4.
In turbulence one often observes $S_{r}^{2}(\mu)\sim r^{\zeta_{2}}$ over an inertial range. Heuristically this corresponds to $s\approx\zeta_{2}/2$, so the coverage term in Corollary 4.11 can be read as a direct proxy for inertial-range scaling.
5.4 Master inequality
Theorem 5.5 (Law-level error: coverage + fit + straightness + discretization).
Fix an effective resolution cutoff $K_{c}$ and assume the structure-function modulus hypothesis used in Corollary 4.11. Let $\hat{\mu}_{1}^{(N)}$ be the law obtained by taking $N$ explicit-Euler steps to integrate the learned rectified ODE $\dot{u}_{\tau}=v_{\theta}(u_{\tau},\tau)$ from $\tau=0$ to $\tau=1$. Assume $v_{\theta}(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$. Then, for constants depending on $L$,
@@ 5.4 Master inequality
Proof sketch.
The coverage term is Corollary˜4.11. The continuous-time mismatch between the ideal rectified flow and the learned flow is controlled by Theorem 4.2 via $\varepsilon_{\mathrm{fit}}$ and $\varepsilon_{\mathrm{curv}}$. The discretization term follows from Corollary 5.2 applied to the learned ODE. ∎
Reading (25) against the experiments.
Equation (25) matches the empirical breakdown seen in Section 6: (i) coverage corresponds to reproducing the high-wavenumber energy content (Figure 13); (ii) fit corresponds to the overall statistical accuracy, reflected in $e_{\mu}$, $e_{\sigma}$, and $W_{1}$ (Table 1); (iii) straightness and discretization are exactly what turn “a good model” into “a fast sampler”: ReFlow reaches the same law-level quality using N=8N\!=\!8 steps, whereas diffusion baselines require $N\!\geq\!128$ (Section 5.6).
@@ 5.5 Implementation Details for PDEs
1:Dataset $\{(u_{0},u_{1})\}$ with $u_{1}=\mathcal{S}(u_{0})$; velocity model $v_{\theta}$; noise schedule $\sigma(\tau)$
@@
where $u_{\tau}$ is a low-frequency representation of $u_{1}$, $u_{0}$ is the initial datum, and $\Gamma(\cdot)$ is an MLP-based time embedding:
@@
Within the U-Net (see SM J.1), we alternate residual blocks with multi‐head attention layers, downsampling to a bottleneck and then upsampling via skip connections. A final $1\times 1$ projection recovers either the field increment $\hat{u}\approx u_{1}-u_{0}$ or a noise residual. By injecting the rectification time $\Gamma(\tau)$ into each block, the model performs a fully deterministic trajectory integration instead of stochastic diffusion.
@@
$\frac{du_{\tau}}{d\tau}=v_{\theta}\bigl(u_{\tau},\,u_{0},\,\tau\bigr),\quad u_{\tau}|_{\tau=0}=\xi,\quad\xi\sim\mathcal{N}(0,I).$
@@
Efficiency and accuracy. Because rectified flows explicitly encourage a “straighter” velocity field, accurate solutions often require fewer integration steps than their diffusion-based counterparts. This advantage can be crucial in computational fluid dynamics and other PDE applications where each solver step is expensive. See SM M.3 and P for a detailed comparison of step-count vs. accuracy trade-offs.
@@ 5.6 Computational Efficiency Comparison
Diffusion surrogates (e.g. GenCFD) generate samples by integrating a reverse-time SDE whose trajectory bends through latent noise; even with accelerated samplers such as DDPM or DPM-solver the procedure still requires tens of score-network evaluations. Rectified flows, in contrast, follow a single deterministic ODE whose nearly straight path tolerates large or adaptive steps, so inference needs only a handful of evaluations. Figure 2 quantifies this gap, showing how the runtime of diffusion models scales almost linearly with the number of steps, whereas rectified flows remain fast even as network size grows.
Figure 2: Inference cost versus model size. Rectified flows need 4488 ODE steps vs. GenCFD’s 6464128128 SDE steps. U/̄Net depth varies from $4$ to 88 blocks (parameters $5.5$ M to $10.2$ M).
@@ 6.1 Experimental Setup
We benchmark our method, termed henceforth as ReFlow, against four neural surrogate models. GenCFD and GenCFD $\circ$ FNO are conditional diffusion surrogates following the framework introduced earlier in this section, whereas FNO and UViT serve as one-shot deterministic baselines. Detailed descriptions of all baselines are provided in SM L.
@@
    Shear Layer in 2D. This setup models the evolution of an incompressible shear layer, initially concentrated near a horizontal interface, which rolls up into intricate vortical structures over time. We generate approximate solutions by numerically solving the 2D Navier–Stokes system (with periodic boundaries) at a resolution of $512^{2}$, then downsample to $128^{2}$ for training and testing. The data has been generated with the spectral viscosity code Tadmor (1989); Rohner and Mishra (2024).
@@
    Cloud-Shock Interaction in 2D. In this popular compressible-flow test, a shock wave moves through a higher-density “cloud” region, triggering complex wave interactions and small-scale turbulent mixing in the wake. We solve the compressible Euler system at an effective resolution of $256^{2}$ using a GPU-optimized finite-volume ALSVINN code of Lye (2020); Fjordholm et al. (2020), then also downsample to $128^{2}$.
@@
    Richtmeyer–Meshkov in 2D. In this case, a shock wave interacts with a density interface separating two fluids, depositing vorticity and triggering vigorous mixing. We numerically solve the compressible Euler system on a $256^{2}$ grid, then downsample to $128^{2}$ to create the training and test data. The data has been generated with a high-resolution finite-volume ALSVINN code of Fjordholm et al. (2020)
@@
being the probability distribution specified in the Supplementary Material (SMK.2). For each $u_{0}$, the PDE solver (e.g. a spectral or finite-volume code) is run forward in time to produce high-fidelity snapshots, which serve as training data.
@@
To create a “ground truth” ensemble for comparison under a fixed $\widehat{u}_{0}$, we employ an ensemble perturbation procedure (detailed in SM  K.2), where $\widehat{u}_{0}$ is slightly perturbed to generate multiple nearby initial states. Each perturbed state is then advanced using a high-resolution PDE solver (e.g. the Azeban code for incompressible flows, or a finite-volume code for compressible problems), yielding a set of flow realizations. This set approximates the conditional measure in question, i.e. $p\!\bigl(u(t)\mid\widehat{u}_{0}\bigr)$.
@@
    Distribution shift. We consider Dirac initial data $\delta_{\widehat{u}_{0}}$ where $\widehat{u}_{0}\sim\nu$, for a different distribution $\nu\neq\mu$. This scenario probes the ability of our model and baselines to generalize under shifts in the initial condition distribution. Specific details on $\nu$ are provided in SM K.2.
We will show results in both regimes to assess how well rectified flows can capture the spread of chaotic solutions starting from fixed initial states–and how it compares to baselines.
@@ 6.2 Evaluation Metrics
We evaluate the generative performance of each method using three main metrics, with details provided in SM M:
@@
One-point Wasserstein as a k=1k\!=\!1 marginal metric.
Let $\mu$ and $\widehat{\mu}$ be laws on fields $u:D\to\mathbb{R}^{m}$. For each $x\in D$ define $\mu_{x}:=(\mathrm{ev}_{x})_{\#}\mu$ and $\widehat{\mu}_{x}:=(\mathrm{ev}_{x})_{\#}\widehat{\mu}$. Our reported metric is
$W_{1}^{(1)}(\mu,\widehat{\mu}):=\int_{D}W_{1}(\mu_{x},\widehat{\mu}_{x})\,dx.$
By Kantorovich–Rubinstein duality, $W_{1}(\mu_{x},\widehat{\mu}_{x})=\sup_{\operatorname{Lip}(\varphi)\leq 1}\big|\int\varphi\,d\mu_{x}-\int\varphi\,d\widehat{\mu}_{x}\big|$, so $W_{1}^{(1)}$ controls the averaged discrepancy of all pointwise 1-Lipschitz observables. This is the precise sense in which Table 1 measures agreement of k=1k\!=\!1 correlation statistics, in line with the statistical-solution viewpoint Lanthaler et al. (2021a).
Table 1: Comparison of models across datasets CS2D, SL2D, and RM2D, in terms of mean relative error ($e_{\mu}$), standard deviation relative error ($e_{\sigma}$), and Wasserstein-1 distance ($W_{1}$). Best-performing model is colored in Blue and second-best performing model in Orange.
@@
Baselines and Evaluation Settings. To stay within reasonable training times for our infrastructure, we restrict the total number of training iterations up to 120,000 for each model and recorded the best achievable results. Under this budget, only a reduced-capacity GenCFD (with two downsampling stages in its U-Net backbone, versus three in Rectified Flow) can match the accuracy reported in Table 1. We also evaluated larger GenCFD variants trained for 500,000 iterations, but observed no meaningful improvement over the results shown here. Meanwhile, at inference time Rectified Flow requires only 8 denoising steps to achieve comparable or superior accuracy, whereas diffusion baselines (including GenCFD) need 128 steps, underscoring RF’s efficiency in both training and sampling. FNO and UViT also require smaller models than ReFlow for optimal performance.
Since FNO, and UViT produce deterministic predictions, we approximate their spread by an ensemble perturbation strategy (SM  K.2), where each baseline is fed slightly perturbed inputs and outputs are aggregated to form a sample set. In contrast, both the original GenCFD (baseline) and our ReFlow are inherently probabilistic, requiring no auxiliary procedure to generate diverse samples.
Figure 3: Energy-spectrum evolution for a random cloud–shock initial datum. ReFlow (blue) tracks high-wavenumber energy faster and more accurately than GenCFD (red).
@@
Moreover as visualized in the SM, ReFlow generates samples of very high quality, matching the qualitative features of the ground truth and capturing features such as (interacting) vortices as well as propagating shock waves very accurately. Another aspect where ReFlow really shines in being able to generate the correct point pdfs. Although already indicated in the very low 1-point Wasserstein distance errors from Table 1, we present 1-point pdfs in the SM to illustrate this observation.
Another key indicator of the quality of approximation of multiscale PDE solutions is the energy spectrum Molinaro et al. (2024). In Figure 13, we plot the (log) spectra for the energy variable of a randomly chosen Cloud Shock initial datum at four different Diffusion times, $\tau\in\{0.25,\,0.50,\,0.75,\,1.00\}$ to observe that i) the solution is indeed highly multiscale, with spectra decaying as a power law over a range of frequencies ii) ReFloW approximates the spectra accurately within a few (5) steps and iii) on the other hand, for these few steps in solving the underlying reverse SDE, a score-based Diffusion model such as GenCFD, struggles to approximate the relevant small scales. It will require much more sampling steps to do so. These results also showcase how ReFlow is able to infer the correct statistical solutions with significantly less ODE solve steps and is hence, much faster, than a diffusion based model such as GenCFD. This gain in inference speed is also reinforced from Figure 4, where we compare ReFlow (RF) and GenCFD reconstructions of the density field $\rho$ for the Cloud-Shock interaction over five diffusion times $\tau\in\{0.00,\,0.25,\,0.50,\,0.75,\,1.00\}$. We see that ReFlow inpaints fine-scale features in far fewer steps, whereas GenCFD remains visibly much more noisy at these small number of steps. See SM M.3 for further visualizations of this effect.
Figure 4: Density trajectories for Sample 1: ReFlow (top) versus GenCFD (bottom).
7 Additional empirical analysis (integrators, robustness, and diversity)
This section consolidates additional experiments conducted in response to reviewer requests, with the goal of making the computational trade-offs and the role of the curvature-aware controller explicit. The emphasis is on (i) solver choices for the learned rectified-flow ODE, (ii) robustness of the controller to its hyperparameters, and (iii) distributional fidelity / diversity under the macro–micro evaluation protocol.
7.1 Protocol recap: macro–micro ensembles, OOD macros, and metrics
Macro–micro conditional ensembles.
To evaluate statistical fidelity rather than only pointwise accuracy, we follow a macro–micro protocol. For each dataset, we select $N_{\mathrm{macro}}$ macroscopic states (“macros”) and for each macro generate $N_{\mathrm{micro}}$ microscopic perturbations (“micros”) yielding a reference ensemble of size $N_{\mathrm{macro}}\times N_{\mathrm{micro}}$. This makes diversity concrete: for each fixed macro, the ground-truth solver produces an empirical conditional distribution over micros, and the model produces its own empirical conditional distribution. We compare these distributions directly via Wasserstein distances (below), which is the relevant question for statistical surrogates.
OOD macro selection.
Unless otherwise stated, we report results in an out-of-distribution (OOD) regime by choosing $N_{\mathrm{macro}}=10$ macros from shifted initial-condition families / parameter ranges, and $N_{\mathrm{micro}}=20$ micros per macro (200 conditional samples per dataset). All models are frozen; only the sampling/integration procedure is varied.
Metrics.
We report: (i) relative $L^{2}$ error against the PDE solution,
$\mathrm{Rel}\text{-}L^{2}(\widehat{u},u):=\frac{\|\widehat{u}-u\|_{L^{2}}}{\|u\|_{L^{2}}},$ (26)
(ii) NFE (neural function evaluations), i.e. the number of velocity-field evaluations per generated sample, and (iii) a combined cost–accuracy scalar
$\mathrm{Cost}\times\mathrm{Err}:=(\mathrm{Rel}\text{-}L^{2})\times(\mathrm{NFE}),$ (27)
which is convenient for summarizing the NFE–error plane (lower is better). Distributional fidelity is measured via empirical $W_{1}$ between conditional ensembles; we report both aggregated and per-variable $W_{1}$ (Section 7.5).
7.2 Integrator study at sampling time: fixed-step baselines vs. curvature-aware control
A recurring concern is whether the gains are merely a consequence of choosing a particular numerical solver. To isolate this, we fix the trained rectified-flow model and vary only the integrator used to solve the rectified-flow ODE at sampling time, starting from noise and integrating from rectified time $\tau=0$ to $\tau=1$. We include a reference implementation from torchdiffeq and standard explicit fixed-step schemes. We report the aggregated results on ShearLayer2D in the OOD macro–micro regime; the second benchmark exhibits the same qualitative trends and is reported in the appendix.